OpenStudy (anonymous):

Find the equation, in standard from, of the line passing through the specified points with the specified slopes.

OpenStudy (anonymous):

Point (5,11) slope of -3

OpenStudy (saifoo.khan):

\[y - y_1 = m(x - x_1)\]

OpenStudy (saifoo.khan):

Insert the values and solve.

OpenStudy (saifoo.khan):

\[y - (1) = -3(x - 5)\]

OpenStudy (anonymous):

Wait so what do i do?

OpenStudy (saifoo.khan):

solve it to something like, Standard form (Ax + By = C)

OpenStudy (anonymous):

\[y-11=-3(x-5)\]

OpenStudy (anonymous):

can u solve from here ?

OpenStudy (saifoo.khan):

oh! i missed "11"

OpenStudy (anonymous):

No i am sorry i am really confused...

OpenStudy (anonymous):

\[y=-3x+15+11\]\[y=-3x+26\]\[3x+y=26\]

OpenStudy (saifoo.khan):